# Quantifying community evolution in temporal networks

### v2

@@ II-D1 Union-NMI
The first item represents the mutual information of the partition of common nodes. The second item represents the mutual information calculated by the virtual community added in $L_{1}$ and all communities in $L_{2}$, which measures the complexity of node flow during the transition from $G_{1}$ to $G_{2}$, and a smaller value represents a more complex flow of nodes in the virtual community, indicating that they have been assigned to multiple different communities during the evolution process. The third item represents the new virtual community in $L_{2}$ and the mutual information of all communities in $L_{1}$, and it measures the complexity of node computations during the transition from $G_{1}$ to $G_{2}$. The mutual information is normalised by the arithmetic average of entropy in two labelling schemes, which are for $i=\{1,2\}$ and $j=2$ if $i=1$ or $j=1$ if $i=2$,
$H_{U}(L_{i})=$ $-\sum_{r\neq m^{(i)}}\frac{n^{(i)}_{r}}{N_{U}}\log\left(\frac{n^{(i)}_{r}}{N_{U}}\right)$ (3)
$-\frac{|V_{i}-V_{j}|}{N_{U}}\log\left(\frac{|V_{i}-V_{j}|}{N_{U}}\right),$
where ||$\cdot$|| ⋅ | refers to the cardinality of a set. The first term represents the entropy of the original community label after introducing the virtual community, and the second term represents the entropy introduced due to the virtual community. Combining Eq. 3 and Eq. 2, the Union-NMI is $\operatorname{UNMI}(L_{1},L_{2})=\frac{-2I_{U}(L_{1};L_{2})}{H_{U}(L_{1})+H_{U}(L_{2})}$ just as it was in Eq. 1.
II-D2 Intersection-NMI
The second proposal is to align the node set by taking the intersection set of the partitions, which evaluates the overlap in information between two community structures. INMI can be interpreted as a measure of coherence or consistency between the two structures. High INMI means that the shared community structure captures a significant portion of the information from both original structures. It emphasises the regions where the structures agree, making it useful for identifying core similarities. Here, the number of common nodes in $G_{1}$ and $G_{2}$ is denoted by $N_{I}=|V_{1}\cap V_{2}|$ and let $q_{r}^{(1)}$ be the number of nodes that are labelled $r$ in $L_{1}$ and are not in $G_{2}$, and $q_{r}^{(2)}$ be the number of nodes that are labelled $r$ in $L_{2}$ and are not in $G_{1}$ (Fig. 1(b)). The mutual information is
$I_{I}(L_{1};L_{2})=\sum_{r,s}\frac{n^{(1,2)}_{r,s}}{N_{I}}\log\left(\frac{n^{(1,2)}_{r,s}N_{I}}{(n^{(1)}_{r}-q_{r}^{(1)})(n^{(2)}_{s}-q_{s}^{(2)})}\right).$
@@ II-D2 Intersection-NMI
$\small H_{I}(L_{i})=-\sum_{r}\frac{n^{(i)}_{r}-q_{r}^{(i)}}{N_{I}}\log\left(\frac{n^{(i)}_{r}-q_{r}^{(i)}}{N_{I}}\right).$ (5)
Combining Eq. 4 and Eq. 5 the normalised mutual information is $\operatorname{INMI}(L_{1},L_{2})=\frac{-2I_{I}(L_{1};L_{2})}{H_{I}(L_{1})+H_{I}(L_{2})}$ again the same equation as before but with different terms.
@@
Because the INMI focuses only on nodes that occur in both $G_{1}$ and $G_{2}$ if the nodes in the intersection $V_{1}\cap V_{2}$ remain in the same community then the INMI will be one irrespective of how many extra nodes are in $G_{1}$ and $G_{2}$. This invariance occurs because INMI is designed to measure the similarity based only on the set of nodes present in both partitions being compared, ignoring any nodes that are unique to one partition. However, the INMI value changes when the nodes’ community labels are shuffled. As the shuffle ratio increases, the community structure similarity, as captured by the INMI, gradually decreases. This decrease is due to the reduction in mutual information and the corresponding increase in entropy, reflecting a loss of community structure coherence among shared nodes.
III RESULTS
@@
In this section, we apply our community structure similarity measure to synthetic temporal networks and multiple empirical data sets. In the synthetic temporal networks, we create a very simple synthetic model to generate communities which change quickly or slowly. We follow this with the investigation of five real-world networks.
III-A Validation on synthetic temporal network
@@
Figure 2: Example dynamic community structures of temporal networks with $n=400$ nodes and their pair-to-pair UNMI, INMI measurement, where $\phi\in[0,1]$ is a parameter which, when high, means nodes leave/arrive in the network quickly (it increases from left to right), and $\psi\in[0,1]$ is a parameter which, when high, means nodes move to new communities quickly (it increases from top to bottom). The first graph for each set of experiments represents the community affiliation of nodes in different snapshots. We use four colours to represent the four communities specified in this experiment ( Refer to caption ). The two subsequent diagrams show the similarity measured by UNMI and INMI in community structure between each pair of network slices. (0 Refer to caption 1).
@@
First, we create a candidate node pool with $N$ nodes, each of which is assigned an initial community label $C_{i}$ chosen at random from equiprobable communities (here we choose four). From these $N$ nodes, $n<N$ are chosen at random to form the initial network. Here we choose $N=500$ and $n=400$. A parameter $\phi\in[0,1]$ gives the probability a node will leave the network and be replaced by another node in the pool to keep the size of the network constant. A higher value of $\phi$ means nodes swap between the network and the pool often (they keep a memory of their community when not in the network). Following this, at each iteration, every one of the nodes now in the network will change their community labels with probability $\psi\in[0,1]$ picking one of the remaining three labels with equal probability. A higher value of $\psi$ means communities are fast-moving and nodes change between them quickly. Because we assign nodes to communities using a model there is no need to use a community detection algorithm and our simulated “network” in fact has no network structure. We consider values of $\phi\in\{0,0.001,0.01,0.1\}$ and values of $\psi\in\{0.001,0.01,0.1\}$ and run the simulation for 50 iterations measuring the UNIM and IMNI for each.
@@
The results are shown in Fig. 2 along with a plot showing the membership of each community (with white showing nodes in the pool, not the network). Moving down the diagram means increasing values of $\phi$ (nodes move in and out of the network more rapidly). Moving rightward on the diagram means nodes move between communities more rapidly. For example, at the top left (Fig. 2(a)) nodes stay in the network and change community very slowly. As expected all values of INMI and UNMI are high (red). Conversely, at the bottom right (Fig. 2(l)) nodes move in and out of the network rapidly and change community quickly and the UNMI and INMI are both low except for the diagonal where the same or similar time periods are compared. The UNMI and INMI measures disagree on how similar communities are when we have a situation where the node set changes rapidly but the community labels stay relatively constant (Fig. 2(j,k)) – here INMI finds that communities are similar because the community labels themselves have not changed much but UNMI finds a smaller degree of similarity because the actual node set could be quite different. We argue that in this case, both methods provide complementary information, neither is superior both are needed to get the full picture. In situations where communities change relatively slowly (Fig. 2(e,h)) we can see that the central band is thicker indicating slowly moving stable communities.
III-B Real-world networks
@@
TABLE I: Summary of datasets. An event is a connection between two nodes at a given time. One or more events between two nodes count as a single edge.
@@
We now test the effectiveness of the proposed Union-NMI (UNMI) and Intersection-NMI (INMI) metrics on several real-world temporal networks (see I). The networks are all temporal with events connecting nodes occurring at well-defined times. In this experiment, we pick a window size for each network using the methods described in the supplemental information. Among them, email-EU-core [28] is a mail interactive network of European research institutions, Math Overflow [28] is an interactive network of online question and answer platform, arXiv Hep-TH [35] is a paper citation network of high energy physics section on arXiv. SubReddit [22] is an interactive network between subreddits (a SubReddit is a community of users) on the Reddit platform, and NFT [24] (non-fungible token) is an online transaction network. In the preprocessing, we removed the self-loops in the networks and trimmed the data where the network was sparse. For details on data sets and data preprocessing, see the supporting information. We use the Raphtory software to create windows of the correct size for each graph and partition nodes into communities using the Louvain algorithm [3]. We slide the window by ten percent of the window size and the process. We then look at the UNMI and INMI for these sliding windows producing a two-dimensional heat map as with the synthetic data.
Figure 3: The (a) edge/node counts and (b) similarity in community structures measured by UNMI and INMI in the email-EU-core network.
Figure 4: The (a) edge/node counts and (b) similarity in community structures measured by UNMI and INMI in the Math Overflow network.
Figure 5: The (a) edge/node counts and (b) similarity in community structures measured by UNMI and INMI in the arXiv HEP-TH network.
Figure 6: The (a) edge/node counts and (b) similarity in community structures measured by UNMI and INMI in the Subreddit network.
Figure 7: The (a) edge/node counts and (b) similarity in community structures measured by UNMI and INMI in the NFT network.
@@
In the email-EU-core dataset (Fig. 3), which captures email communications within a large European research institution, we observed a significant decrease in UNMI readings around December 2003 and 2004, corresponding to the Christmas holiday periods. During the same time, INMI readings showed a notable increase but the UNMI had a decrease (see zoom in). This suggests that while there were substantial changes in the overall network—reflected by the varying node sets due to personnel changes and reduced communication activity—the core group of staff who remained active maintained similar interaction patterns. This shows how the measures respond differently to a change in network membership and is consistent with the results on synthetic data. The INMI is consistently higher because in this network the actual participants change greatly between windows as different members of the organisation become more or less active in sending and receiving emails. The network size remains constant over time and the thickness of the diagonal shows that communities are being formed at the same rate throughout.
@@
For the Math Overflow dataset (Fig. 4), an online platform where users ask and answer mathematical questions, the results showed that communities are slowly forming but have not yet established stable structures. The open and fluid interaction model of the platform allows users to freely engage in discussions without strict group boundaries, making it challenging for strong communities to form. Although users change, both INMI and UNMI show the main diagonal thickening over time indicating that as time. This could indicate that as time continues users begin to form loosely connected communities, leading to a slow but steady strengthening of community similarity among the core users. The graph here shows the whole history of the network so we might expect no coherent communities at the beginning of the data set.
@@
In the HEP-TH citation network (Fig. 5), representing high-energy physics theory papers from the arXiv repository. The initial part of this data has been removed as the number of papers per year was very small and we can see the rate of nodes being added increases considerably. UNMI and INMI both begin high and in particular at the beginning of the data set, where there are few papers, the INMI is quite high. We can see the diagonal of the plot gets thinner for both as time goes on, perhaps indicating that as the number of contributors grows larger and larger that core forming a community is smaller by comparison. This reflects the dynamic nature of academic research, where the number of publications increases as time continues and evolving interests and expanding collaborations can fragment initially cohesive communities.
@@
In the NFT transaction network (Fig. 7), which records buying and selling activities in the NFT market, we observed that during the last year of the dataset, INMI readings were significantly high, while UNMI decreased. The decrease in UNMI suggests that a large number of new participants entered the market, increasing the node count and altering the overall network structure, thus reducing similarity when considering all nodes. However, the high INMI readings indicate that core buyers continued to interact predominantly with trusted users, maintaining consistent trading relationships. This stability among persistent participants led to more similar community structures within the core network. The findings align with previous research [24] suggesting that despite market expansion, the NFT ecosystem is driven by a small number of active participants who engage in frequent transactions within established communities.
IV DISCUSSION

### v3

@@ 3.2.1 Union-AMI
The first term represents the mutual information of the partition of common nodes. The second term represents the mutual information calculated by the virtual community added in $L_{1}$ and all communities in $L_{2}$, which measures the complexity of node flow during the transition from $G_{1}$ to $G_{2}$, and a smaller value represents a more complex flow of nodes in the virtual community, indicating that they have been assigned to multiple different communities during the evolution process. The third term represents the new virtual community in $L_{2}$ and the mutual information of all communities in $L_{1}$, and it measures the complexity of node computations during the transition from $G_{1}$ to $G_{2}$. The mutual information is normalised by the arithmetic average of entropy in two labelling schemes, which are for $i=\{1,2\}$ and $j=2$ if $i=1$ or $j=1$ if $i=2$,
$H_{U}(L_{i})=-\sum_{r\neq m^{(i)}}\frac{n^{(i)}_{r}}{N_{U}}\log\left(\frac{n^{(i)}_{r}}{N_{U}}\right)-\frac{|V_{i}-V_{j}|}{N_{U}}\log\left(\frac{|V_{i}-V_{j}|}{N_{U}}\right),$ (3)
where $|\cdot|$ refers to the cardinality of a set. The first term represents the entropy of the original community label after introducing the virtual community, and the second term represents the entropy introduced due to the virtual community. Subsequently, by subtracting the average value of the mutual information that can be generated by random partitioning under the null model assumption. We obtained UAMI. The expression of the expectation term is as follows:
$\mathbb{E}_{U}[\operatorname{MI}]=\sum_{r=L_{1}\cup\{m^{(1)}\}}\sum_{s=L_{2}\cup\{m^{(2)}\}}\sum_{n=n_{r,s}^{\min}}^{n_{r,s}^{\max}}\frac{\binom{n_{r}^{(1)}}{n}\binom{N_{U}-n_{r}^{(1)}}{n_{s}^{(2)}-n}}{\binom{N_{U}}{n_{s}^{(2)}}}\times\frac{n}{N_{U}}\log\left(\frac{nN_{U}}{n_{r}^{(1)}n_{s}^{(2)}}\right).$ (4)
Combining Eq. 2 and Eq. 3, then the expression of UAMI is $\frac{2(I_{I}(L_{1};L_{2})-\mathbb{E}_{U}[\operatorname{MI}])}{H_{I}(L_{1})+H_{I}(L_{2})-\mathbb{E}_{U}[\operatorname{MI}]}$.
3.2.2 Intersection-AMI
The second proposal is to align the node set by taking the intersection set of the partitions, which evaluates the overlap in information between two community structures. The intersection extension can be interpreted as a measure of coherence or consistency between the two structures. A higher value in the intersection extension means that the shared community structure captures a significant portion of the information from both original structures. It emphasises the regions where the structures agree, making it useful for identifying core similarities. Here, the number of common nodes in $G_{1}$ and $G_{2}$ is denoted by $N_{I}=|V_{1}\cap V_{2}|$ and let $q_{r}^{(1)}$ be the number of nodes that are labelled $r$ in $L_{1}$ and are also in $G_{2}$, and $q_{s}^{(2)}$ be the number of nodes that are labelled $r$ in $L_{2}$ and are also in $G_{1}$ and $q_{r,s}^{(1,2)}$ is the number of nodes labelled $r$ in $L_{1}$ and $s$ in $L_{2}$ and they appears in both $G_{1}$ and $G_{2}$ (Fig. 1(b)). The mutual information is
$I_{I}(L_{1};L_{2})=\sum_{r,s}\frac{q^{(1,2)}_{r,s}}{N_{I}}\log\left(\frac{q^{(1,2)}_{r,s}N_{I}}{q_{r}^{(1)}q_{s}^{(2)}}\right).$ (5)
@@ 3.2.2 Intersection-AMI
$H_{I}(L_{i})=-\sum_{r}\frac{q_{r}^{(i)}}{N_{I}}\log\left(\frac{q_{r}^{(i)}}{N_{I}}\right).$ (6)
Again, we subtract the expectation to eliminate the bias caused by the number of communities. When taking the intersection, the expectation term is expressed as:
$\mathbb{E}_{I}[\mathrm{MI}]=\sum_{r}\sum_{s}\sum_{q=q_{r,s}^{\min}}^{q_{r,s}^{\max}}\frac{\binom{q_{r}^{(1)}}{q}\binom{N_{I}-q_{r}^{(2)}}{q_{s}^{(2)}-q}}{\binom{N_{I}}{q_{s}^{(2)}}}\times\frac{q}{N_{I}}\log\left(\frac{qN_{I}}{q_{r}^{(1)}q_{s}^{(2)}}\right).$ (7)
Then the expression of IAMI is $\frac{2(I_{I}(L_{1};L_{2})-\mathbb{E}_{I}[\operatorname{MI}])}{H_{I}(L_{1})+H_{I}(L_{2})-\mathbb{E}_{I}[\operatorname{MI}]}$.
@@
Because the intersection extension focuses only on nodes that occur in both $G_{1}$ and $G_{2}$ if the nodes in the intersection $V_{1}\cap V_{2}$ remain in the same community then the intersection will be one irrespective of how many extra nodes are in $G_{1}$ and $G_{2}$. This invariance occurs because the IAMI is designed to measure the similarity based only on the set of nodes present in both partitions being compared, ignoring any nodes that are unique to one partition. However, the IAMI value changes when the nodes’ community labels are shuffled. As the shuffle ratio increases, the community structure similarity, as captured by the IAMI, gradually decreases. This decrease is due to the reduction in mutual information and the corresponding increase in entropy, reflecting a loss of community structure coherence among shared nodes.
4 Results
@@
In this section, we apply our community structure similarity measure to synthetic temporal networks and multiple empirical data sets. In the synthetic temporal networks, we create a very simple synthetic model to generate communities which change quickly or slowly. We follow this with the investigation of five real-world networks. In addition, we replicated the methods of the relevant studies [wang2022quantifying], demonstrating the reliability of our approach and enabling faster computations on the datasets (please refer to supplementary Fig. 9.
4.1 Validation on synthetic data
@@
Figure 2: Example dynamic community structures with $n=400$ nodes and their pair-to-pair UAMI, IAMI measurement, where $p_{\mathrm{rot}}\in[0,1]$ is a parameter which, when high, means nodes leave/arrive in the network quickly (it increases from left to right), and $p_{\mathrm{move}}\in[0,1]$ is a parameter which, when high, means nodes move to new communities quickly (it increases from top to bottom). The first graph for each set of experiments represents the community assignments of nodes in different snapshots. We use four colours to represent the four communities specified in this experiment ( Refer to caption ). The two subsequent diagrams show the similarity measured by UAMI and IAMI in community structure between each pair of network slices. (0 Refer to caption 1).
@@
First, we create a candidate node pool with $N$ nodes, each of which is assigned an initial community label $C_{i}$ chosen at random from equiprobable communities (here we choose four). From these $N$ nodes, $n<N$ are chosen at random to form the initial network. Here we choose $N=500$ and $n=400$. A parameter $p_{\mathrm{rot}}\in[0,1]$ gives the probability a node will leave the network and be replaced by another node in the pool to keep the size of the network constant. A higher value of $p_{\mathrm{rot}}$ means nodes swap between the network and the pool often (they keep a memory of their community when not in the network). Following this, at each iteration, every one of the nodes now in the network will change their community labels with probability $p_{\mathrm{move}}\in[0,1]$ picking one of the remaining three labels with equal probability. A higher value of $p_{\mathrm{move}}$ means communities are fast-moving and nodes change between them quickly. We consider values of $p_{\mathrm{rot}}\in\{0,0.001,0.01,0.03,0.1,0.3\}$ and values of $p_{\mathrm{move}}\in\{0,0.001,0.01,0.03,0.1,0.3\}$ and run the simulation for 50 iterations measuring the UAMI and IAMI for each. In Fig. 2, we present the experimental results of UAMI and IAMI under representative $p_{\mathrm{rot}}$ and $p_{\mathrm{move}}$. For more experimental results at different values, please refer to the supplementary Fig. 6-8.
@@
The results are shown in Fig. 2 along with a plot showing the membership of each community (with white showing nodes in the pool, not the network). Moving down the diagram means increasing values of $p_{\mathrm{rot}}$ (nodes move in and out of the network more rapidly). Moving rightward on the diagram means nodes move between communities more rapidly. For example, at the top left (Fig. 2(a)) nodes stay in the network and change community very slowly. As expected all values of IAMI and UAMI are high (red). Conversely, at the bottom right (Fig. 2(l)) nodes move in and out of the network rapidly and change community quickly and the UAMI and IAMI are both low except for the diagonal where the same or similar time periods are compared. The UAMI and IAMI measures degree on how similar communities are when we have a situation where the node set changes rapidly but the community labels stay relatively constant (Fig. 2(j,k)) – here IAMI finds that communities are similar because the community labels themselves have not changed much but UAMI finds a smaller degree of similarity because the actual node set could be quite different. We argue that in this case, both methods provide complementary information, neither is superior both are needed to get the full picture. In situations where communities change relatively slowly (Fig. 2(e,h)) we can see that the central band is thicker indicating slowly moving communities.
@@
Table 1: Summary of datasets. An event is a connection between two nodes at a given time. One or more events between two nodes count as a single edge.
4.2 Real-world networks
@@
We now test the effectiveness of the proposed Union-AMI (UAMI) and Intersection-AMI (IAMI) metrics on several real-world temporal networks (See Tab. 1). The networks are all temporal with events connecting nodes occurring at well-defined times. In this experiment, we pick a window size for each network using the methods described in the supplemental information. Among them, email-EU-core [paranjape2017motifs] is a mail interactive network of European research institutions, Math Overflow [paranjape2017motifs] is an interactive network of online question and answer platform, arXiv Hep-TH [ryan2015the] is a paper citation network of high energy physics section on arXiv. SubReddit [kumar2018community] is an interactive network between subreddits (a SubReddit is a community of users) on the Reddit platform, and NFT [nadini2021mapping] (non-fungible token) is an online transaction network. In the preprocessing, we removed the self-loops in the networks and trimmed the data where the network was sparse. For details on data sets and data preprocessing, see the supporting information. We use the Raphtory software [steer2023raphtory] to create windows of the correct size for each graph and partition nodes into communities using the Louvain algorithm [blondel2008fast]. For the relevant statistical results, please refer to supplementary Fig. 1-5 in the supporting information. We slide the window by ten percent of the window size and the process. We then look at the UAMI and IAMI for these sliding windows producing a two-dimensional heat map as with the synthetic data.
Figure 3: (a) shows edge and node counts. (b) shows the proportion of nodes appears only in the $1^{\text{st}}$ window, the $2^{\text{nd}}$ window, and appears in both windows. (c) shows the similarity in community measures in the email-EU-core network.
Figure 4: (a) shows edge and node counts. (b) shows the proportion of nodes appears only in the $1^{\text{st}}$ window, the $2^{\text{nd}}$ window, and appears in both windows. (c) shows the similarity in community measures in the Math Overflow network.
Figure 5: (a) shows edge and node counts. (b) shows the proportion of nodes appears only in the $1^{\text{st}}$ window, the $2^{\text{nd}}$ window, and appears in both windows. (c) shows the similarity in community measures in the arXiv HEP-TH network.
Figure 6: (a) shows edge and node counts. (b) shows the proportion of nodes appears only in the $1^{\text{st}}$ window, the $2^{\text{nd}}$ window, and appears in both windows. (c) shows the similarity in community measures in the subreddit network.
Figure 7: (a) shows edge and node counts. (b) shows the proportion of nodes appears only in the $1^{\text{st}}$ window, the $2^{\text{nd}}$ window, and appears in both windows. (c) shows the similarity in community measures in the NFT trade network.
@@
In the email-EU-core dataset (Fig. 3), which captures email communications within a large European research institution, we observed a significant decrease in UAMI readings around December 2003 and 2004, corresponding to the Christmas holiday periods. During the same time, IAMI readings showed a notable increase but the UAMI had a decrease (see zoom in). This suggests that while there were substantial changes in the overall network—reflected by the varying node sets due to personnel changes and reduced communication activity—the core group of staff who remained active maintained similar interaction patterns. This shows how the measures respond differently to a change in network membership and is consistent with the results on synthetic data. The IAMI is consistently higher because in this network the actual participants change greatly between windows as different members of the organisation become more or less active in sending and receiving emails. The network size remains constant over time and the thickness of the diagonal shows that communities are being formed at the same rate throughout.
@@
For the Math Overflow dataset (Fig. 4), an online platform where users ask and answer mathematical questions, the results showed that communities are slowly forming but have not yet established stable structures. The open and fluid interaction model of the platform allows users to freely engage in discussions without strict group boundaries, making it challenging for strong communities to form. Although users change, both IAMI and UAMI show the main diagonal thickening over time indicating that as time continues users begin to form loosely connected communities. This leads to a slow but steady strengthening of community similarity among the core users. The graph here shows the whole history of the network so we might expect no coherent communities at the beginning of the data set.
@@
In the HEP-TH citation network (Fig. 5), representing high-energy physics theory papers from the arXiv repository. The initial part of this data has been removed as the number of papers per year was very small and we can see the rate of nodes being added increases considerably. UAMI and IAMI both begin high and in particular at the beginning of the data set, where there are few papers, the IAMI is quite high. We can see the diagonal of the plot gets thinner for both as time goes on, perhaps indicating that as the number of contributors grows larger and larger that core forming a community is smaller by comparison. This reflects the dynamic nature of academic research, where the number of publications increases as time continues and evolving interests and expanding collaborations can fragment initially cohesive communities.
@@
In the NFT transaction network (Fig. 7), which records buying and selling activities in the NFT market, we observed that during the last year of the dataset, IAMI readings were significantly high, while UAMI decreased. The decrease in UAMI suggests that a large number of new participants entered the market, increasing the node count and altering the overall network structure, thus reducing similarity when considering all nodes. However, the high IAMI readings indicate that core buyers continued to interact predominantly with trusted users, maintaining consistent trading relationships. This stability among persistent participants led to more similar community structures within the core network. The findings align with previous research [nadini2021mapping] suggesting that despite market expansion, the NFT ecosystem is driven by a small number of active participants who engage in frequent transactions within established communities.
5 Discussion